# KVQuant: Towards 10 Million Context Length LLM Inference with KV Cache Quantization

### v3

@@ Appendix D Per-Channel Key Quantization Ablations
int3 PC PC 8.19 12.0
int3 PC PT 6.99 12.0
Appendix E Pre-RoPE Key Quantization Ablations
As shown in Table 8, pre-RoPE Key quantization achieves higher accuracy than post-RoPE quantization, with an improvement of 0.65 perplexity for 3-bit quantization with the LLaMA-7B model. These results show that the rotary positional embeddings make Key quantization more challenging due to mixing pairs of channels with different magnitudes. Pre-RoPE quantization thereby allows for more accurate quantization at low precision.
@@ Appendix E Pre-RoPE Key Quantization Ablations
int3 post-RoPE 6.99 12.0
int3 pre-RoPE 6.34 12.0
Appendix F Sensitivity-Weighted Non-Uniform Quantization Ablations
Table 9 shows perplexity evaluation results across different LLaMA, LLaMA-2, and Mistral models on Wikitext-2 using nf3, nuq3, and nuq3 without using sensitivity-weighting. These results demonstrate the benefits of our sensitivity-weighted non-uniform quantization approach, relative to NormalFloat quantization [6], as we achieve consistent accuracy improvements of up to 0.32 perplexity across different models (with particularly pronounced improvements for larger models). For 3-bit quantization with the LLaMA-7B model, we observe a 0.33 perplexity improvement relative to uniform quantization. The gains relative to uniform quantization are particularly noticeable for 3-bit and 2-bit quantization, where the benefits of non-uniform quantization are more pronounced due to the reduced precision. These results also demonstrate the necessity of our sensitivity-weighting approach in order to derive performant non-uniform datatypes using a k-means based approach.
Figure 6 shows the relationship between the Fisher information and the normalized activation values for the LLaMA-7B model. For the Value matrices, we observe that the average Fisher information is higher close to the middle, which leads to quantization centroids being pulled closer to the center of the distribution. For the Key matrices, we observe that the average sensitivity across different magnitudes is relatively constant, which leads to wider spacing for the centroids. These results show that using Fisher information to derive a sensitivity-weighted per-layer datatype allows for better representation of Keys and Values.
Figure 6. Relationship between average Fisher information and normalized activation values for the 3-bit quantized LLaMA-7B model with 1% outliers, shown for layer 29. a) Fisher information versus normalized activation values for Keys (with 1% outliers). b) Fisher information versus normalized activation values for Values (with 1% outliers). These results demonstrate how sensitivity-weighted k-means shifts the quantization signposts inward relative to NormalFloat (with more condensed signposts across a slightly narrower range). Additionally, for the Value matrix, the values close to 0 have a higher average sensitivity. This leads to more condensed signpost assignment relative to the Key matrix, demonstrating how our non-uniform approach derives accurate datatypes for layers with differing sensitivity characteristics.
Table 9. Ablation Study: Ablation of our sensitivity-weighted non-uniform datatype for different models on Wikitext-2. All experiments use pre-RoPE per-channel quantization for Keys and per-token quantization for Values (meaning that all configurations are the same as in KVQuant, except for the datatype). We compare against both uniform (int3) and non-uniform (nf3) [6] approaches, as well as with using “unweighted” k-means (i.e., not sensitivity-weighted) to compute the non-uniform quantization signposts. Note that there is slight variation in average bitwidth across models due to the differing hidden dimensions. Results report perplexity with 2K/4K/8K sequence length for LLaMA, LLaMA-2, and Mistral, respectively.
Method LLaMA-7b LLaMA-13b LLaMA-30b LLaMA-65b LLaMA-2-7b LLaMA-2-13b LLaMA-2-70b Mistral-7b Avg. Num. Bits
@@ Appendix F Sensitivity-Weighted Non-Uniform Quantization Ablations
nuq3 6.01 5.34 4.41 3.74 5.49 4.83 3.26 5.03 3.00-3.02
Appendix G Per-Vector Dense-and-Sparse Quantization Ablations
Table 10 shows the performance improvements we observe when isolating a small portion of outliers and storing them in a sparse format. We provide results both with using a single per-matrix outlier threshold, as well as with applying separate outlier thresholds per-vector. In particular, we see greater improvements by employing outlier detection with a different threshold per-channel for Keys and per-token for Values. This provides additional benefits since some values which would be considered outliers for the entire matrix are not actually outliers within a particular channel (so they are not hard to quantize). It is therefore better to directly target the outliers that will skew the quantization range for a particular channel. By removing 1% of outliers using per-vector thresholds, we can achieve an additional 0.25 reduction in perplexity for the LLaMA-7b model at 3 bits, thereby enabling 3-bit quantization with under 0.1 degradation in perplexity.
Table 10. Ablation Study: Perplexity comparison of different outlier isolation methods for LLaMA-7B on Wikitext-2. Per-vector outlier detection allows for significant accuracy improvements relative to per-tensor outlier detection. All nuq4 and nuq3 experiments use per-token quantization for Values and per-channel quantization for Keys (pre-RoPE). “PV” refers to using per-vector outlier thresholds, and “PM” refers to using a single per-matrix outlier threshold.
@@ Appendix G Per-Vector Dense-and-Sparse Quantization Ablations
nuq3 - - 6.01 12.0
nuq3 0.1% PM 5.94 12.2
nuq3 0.1% PV 5.86 12.2
nuq3 1% PM 5.86 13.3
nuq3 1% PV 5.76 13.3
Appendix H Q-Norm Ablations
Table 11 provides perplexity results for LLaMA-7B and LLaMA-13B with and without Q-Norm. As shown in the table, Q-Norm provides noticeable accuracy improvements for 2-bit quantization, but it doesn’t provide significant accuracy improvements for 3-bit or 4-bit quantization. Figure 7 demonstrates how the minimization of distribution shift provided by Q-Norm leads to reduced quantization error (particularly at later layers in the network).
Additionally, we experimented with using per-vector Q-Norm, where we normalize each channel for Keys and each token for Values to ensure the distribution has the same mean and standard deviation post-quantization. For Keys, per-vector Q-Norm requires offline computation of the required normalization per-channel; and for Values, per-vector Q-Norm requires online computation of the required normalization per-token. With per-vector Q-Norm, the normalization parameters must also be stored separately from the per-vector outlier thresholds and cannot be fused with the NUQ datatype. Table 11 also compares employing per-vector Q-Norm with per-matrix Q-Norm. Although we observe similar accuracy with per-vector Q-Norm when incorporating dense-and-sparse quantization, we observe worsened performance when we aren’t using dense-and-sparse quantization. We attribute this to the larger relative shift when applying normalization per-vector, as well as the impacts that significant changes to the centroids can have on outlier values (in the case where we aren’t using dense-and-sparse quantization, large perturbations in these outlier values can lead to significant accuracy loss). The need to only slightly tweak normalization parameters when combating distribution shift is similar to [21], which describes how weight quantization can be improved by slightly adjusting layernorm parameters to avoid distribution shift. Due to the improved performance of per-matrix Q-Norm for dense-only quantization (as well as potential inference overheads with per-vector Q-Norm from having to separately rescale the non-uniform centroids per-vector and compute normalization statistics on-the-fly for Values), we focus on per-matrix Q-Norm. However, there is potential to improve accuracy through fine-grained normalization in future work; for example,  Figure 8 demonstrates how per-vector Q-Norm provides improved performance for the LLaMA-2-70B-32K model when evaluating perplexity on long context lengths.
Table 11. Ablation Study: Perplexity on Wikitext-2 with LLaMA-7B and LLaMA-13B, with and without Q-Norm (including results for both per-matrix (“PM”) and per-vector (“PV”) Q-Norm). For 4-bit and 3-bit quantization, Q-Norm provides minimal perplexity improvements; however, at 2-bit quantization, Q-Norm improves performance by reducing distribution shift. Per-vector Q-Norm performs similarly to per-matrix Q-Norm when incorporating dense-and-sparse quantization, and performs worse than per-matrix Q-Norm without dense-and-sparse quantization. Note that there is slight variation in average bitwidth across models due to the differing hidden dimensions.
Datatype Norm LLaMA-7B LLaMA-13B Avg. Bits (KV Cache)
fp16 - 5.68 5.09 16
nuq4 - 5.73 5.15 4.00
nuq4 PV 5.76 5.16 4.01
nuq4 PM 5.74 5.14 4.00
nuq4-1% - 5.70 5.11 4.32-4.33
nuq4-1% PV 5.70 5.10 4.33
nuq4-1% PM 5.70 5.10 4.32-4.33
nuq3 - 6.01 5.34 3.00
nuq3 PV 6.11 5.42 3.01
nuq3 PM 5.97 5.34 3.00
nuq3-1% - 5.76 5.15 3.32-3.33
nuq3-1% PV 5.75 5.16 3.33
nuq3-1% PM 5.75 5.16 3.32-3.33
nuq2 - 8.70 7.50 2.00
nuq2 PV 34.81 18.56 2.01
nuq2 PM 8.17 7.29 2.00
nuq2-1% - 6.24 5.50 2.32-2.33
nuq2-1% PV 6.06 5.41 2.33
nuq2-1% PM 6.06 5.40 2.32-2.33
@@ Appendix I Attention Sink-Aware Quantization Ablations
Table 12 provides perplexity with and without Attention Sink-Aware quantization across different LLaMA models ranging from 7B to 65B. For all bit widths (4, 3, and 2-bit) and in both dense-only and dense-and-sparse quantization settings, Attention Sink-Aware quantization consistently yields perplexity improvement. Notably, the perplexity gain is more pronounced at lower bit widths and without sparsity.
Table 12. Ablation Study: Perplexity on Wikitext-2 with LLaMA-7B, 13B, 30B, and 65B, with and without Attention Sink-Aware quantization. Attention Sink-Aware quantization consistently improves perplexity across different models and bit widths, particularly with lower bit widths and without sparsity.
Datatype Attention Sink-Aware LLaMA-7B LLaMA-13B LLaMA-30B LLaMA-65B
fp16 - 5.68 5.09 4.10 3.53
nuq4 X 5.73 5.15 4.16 3.57
nuq4 O 5.72 5.13 4.14 3.56
nuq4-1% X 5.70 5.11 4.12 3.54
nuq4-1% O 5.69 5.10 4.11 3.54
nuq3 X 6.01 5.34 4.41 3.74
nuq3 O 5.89 5.25 4.30 3.63
nuq3-1% X 5.76 5.15 4.17 3.59
nuq3-1% O 5.75 5.15 4.16 3.57
nuq2 X 8.17 7.29 7.05 27.17
nuq2 O 7.15 5.81 4.84 4.38
nuq2-1% X 6.06 5.40 4.43 3.76
nuq2-1% O 6.02 5.37 4.36 3.72
Figure 7. Quantization error for 2-bit quantization with and without Q-Norm for the LLaMA-7B model, evaluated on a sample of length 2K from Wikitext-2. a) Quantization error for Keys (with 0% outliers). b) Quantization error for Values (with 0% outliers). c) Quantization error for Keys (with 1% outliers). d) Quantization error for Values (with 1% outliers). As one can see, using Q-Norm is helpful with/without dense-and-sparse quantization and results in improved quantization error, especially for later layers. See Table 11 for the corresponding perplexity results which clearly show the benefit of using Q-Norm.
@@
Table 13 shows accuracy results when using offline calibration for computing the scaling factors for the Keys. For 4-bit quantization, we observe no accuracy loss when calibrating scaling factors offline. For 3-bit quantization, we observe minor accuracy degradation when not employing outlier extraction methods. However, if we remove a small percentage of outliers, then the accuracy with offline calibration is the same as computing the scaling factors online per-channel during evaluation. This demonstrates that when incorporating outlier extraction methods, we are better able to perform offline calibration due to reduced sensitivity to outliers (either to outliers during calibration that exaggerate the quantization range, or to outliers during evaluation that cannot be represented accurately if there weren’t large outliers observed during calibration).
Table 14 shows the runtime for the $topk$ operation for the LLaMA-7B model (which is required for computing outlier thresholds online). It compares the runtime of the $topk$ operation with the runtime for the QKV projections, finding that the $topk$ runtime is 60% of the matrix-vector operation runtime for a single projection layer. The $topk$ operation can also be performed efficiently on the CPU, so we can actually run this operation in parallel with the subsequent linear layer matrix-vector operations on the GPU (which is possible by computing the Value projection before the Key and Query projections). This allows us to compress the activations dynamically without added runtime overhead, thereby enabling online scaling factor computation for the Value tensors.
Table 15 shows the perplexity of the LLaMA-7B model using different numbers of samples during calibration. The results show that perplexity is similar across the range of the number of samples tested for each bit width. This shows that the calibration step does not require a large number of calibration samples to attain high accuracy. Additionally, Table 16 shows how both Fisher information computation and calibration (including k-means) per-layer take only a few minutes for the LLaMA-65B model on a typical server machine. Even if we perform calibration sequentially for each layer, the entire calibration process would take a maximum of 6 hours for the LLaMA-65B model at 4-bit precision.
Table 13. Ablation Study: Model accuracy when using offline calibration for Keys with LLaMA-7B. When incorporating outlier detection, offline calibration for Keys is able to perform comparably with online calibration. All nuq4 and nuq3 experiments use per-token quantization for Values and per-channel quantization for Keys (pre-RoPE), and experiments with outliers use per-vector outlier detection.
@@
nuq4 - 5.73 5.73
nuq4 1% 5.70 5.70
nuq3 - 5.96 6.01
nuq3 1% 5.77 5.76
Table 14. $topk$ runtime on a vector of length 4096 for computing outlier thresholds when using 1% sparsity (compared with the runtime for the QKV matrix multiplications, which are 4096$\times$4096 by 4096 matrix-vector multiplications for the LLaMA-7B model). The runtime is reported on a system with an A6000 GPU and an Intel Xeon Gold 6126 CPU. We find that the runtime for the $topk$ operation is only 60% of the runtime of each matvec operation. Additionally, the $topk$ operation can be performed efficiently on the CPU; we can therefore run this operation in parallel with subsequent linear layer operations on the GPU to compress the activations dynamically without added overhead. Note that the CPU runtime includes the time for copying the vector to the CPU.
@@
QKV Projection GPU - 0.308
$topk$ GPU 1% 0.073
$topk$ CPU 1% 0.059
QKV Projection / $topk$ (Fused) GPU / CPU 1% 0.309
Table 15. Ablation Study: Perplexity on Wikitext-2 for the LLaMA-7B model when using different number of samples of length 2K during calibration.
Method Number of Samples
1 2 4 8 16 32 64
nuq4 5.75 5.74 5.73 5.75 5.73 5.74 5.74
nuq4-1% 5.69 5.70 5.70 5.70 5.70 5.70 5.70
nuq3 6.00 6.00 5.95 5.98 6.01 6.02 5.98
nuq3-1% 5.75 5.77 5.76 5.76 5.76 5.77 5.76
nuq2 8.34 8.32 8.26 8.70 8.61 8.93 8.45
nuq2-1% 6.20 6.22 6.23 6.23 6.24 6.22 6.25
@@
For our empirical evaluation, we use 16 calibration samples of sequence length 2K from the Wikitext-2 training set (as well as the corresponding gradients) to derive the per-channel scaling factors and zero-points, to derive the non-uniform datatypes for both Keys and Values, and to estimate layer-wise sensitivity for mixed-precision experiments. We measured perplexity on both Wikitext-2 and on C4 using a sequence length equal to the maximum context length of the model (2K for LLaMA, 4K for LLaMA-2, and 8K for Mistral-7B). For baseline experiments, we use post-RoPE quantization, both since this is required from an efficiency perspective without a dedicated kernel implementation, and because it provides better accuracy when quantizing Keys per-token as shown in Appendix N.
We make several assumptions in order to estimate average bit widths and KV cache sizes for different approaches. We compute these estimates assuming a sequence length of 128K. For integer quantization, we assume a low-precision integer offset and a 16-bit scaling factor, whereas for NormalFloat and NUQ we assume that the zero-point and offset are each 16-bit. For the sparse matrices, 32-bit integers are assumed for the per-token indices (since we need to support long sequence lengths), and the elements and per-element indices are assumed to be 16-bit. This means that for CSR, the rows are assumed to be 32-bit and the columns and values are assumed to be 16-bit, whereas for CSC, the columns are assumed to be 32-bit and the rows and values are assumed to be 16-bit.
@@
Tables 17 and 18 show perplexity across LLaMA, LLaMA-2, and Mistral models on Wikitext-2 and C4, respectively, using different datatypes. The tables highlight that the NUQ datatype generally outperforms both uniformly quantized INT datatype and non-uniformly quantized NF datatype even after they are incorporated with grouping at the expense of increased bit-widths. This performance can be further improved by extracting 0.1% to 1.0% of outliers. Note that NUQ with a sparsity level 1.0% has a similar memory requirement as INT with a group size of 64, but achieves significant perplexity improvement across all models and bit widths.
Appendix M Full Perplexity Evaluation and MMLU Evaluation
Table 19 shows perplexity evaluation across all models in LLaMA, LLaMA-2, and Mistral models on Wikitext-2. This result demonstrates the benefits of our approach for KV cache compression across different models and bit widths. Table 20 also provides zero-shot evaluation results on MMLU for LLaMA-7B and LLaMA-13B, demonstrating how we can maintain accuracy even for 3-bit KV cache quantization [15]. We used the Language Model Evaluation Harness to run zero-shot evaluation across all MMLU tasks [11].

### v4

@@ Appendix E Per-Channel Key Quantization Ablations
int3 PC PC 223 12.0
int3 PC PT 7.05 12.0
Appendix F Pre-RoPE Key Quantization Ablations
As shown in Table 8, pre-RoPE Key quantization achieves higher accuracy than post-RoPE quantization, with an improvement of 0.82 perplexity for 3-bit quantization with the LLaMA-7B model. These results show that the rotary positional embeddings make Key quantization more challenging due to mixing pairs of channels with different magnitudes. Pre-RoPE quantization thereby allows for more accurate quantization at low precision.
@@ Appendix F Pre-RoPE Key Quantization Ablations
int3 post-RoPE 7.05 12.0
int3 pre-RoPE 6.23 12.0
Appendix G Sensitivity-Weighted Non-Uniform Quantization Ablations
Table 9 shows perplexity evaluation results across different LLaMA, Llama-2, and Mistral models on Wikitext-2 for different datatypes, including nuq3, nuq3 without using sensitivity-weighting, as well as nuq3 with sensitivity-weighting but without accounting for the per-channel scaling factors when performing k-means. We observe particularly noticeable gains relative to uniform quantization for 3-bit and 2-bit quantization, where the benefits of non-uniform quantization are more pronounced due to the reduced precision. These results demonstrate the benefits of our sensitivity-weighted non-uniform quantization approach relative to NormalFloat quantization [8], as we achieve consistent accuracy improvements of up to 0.33 perplexity across different models. These results also demonstrate the necessity of our sensitivity-weighting approach in order to derive performant non-uniform datatypes using a k-means based approach. Additionally, we observe distinct benefits when also accounting for the per-channel scaling factors when performing k-means.
Table 9. Ablation Study: Ablation of our sensitivity-weighted non-uniform datatype for different models on Wikitext-2. All experiments use pre-RoPE per-channel quantization for Keys and per-token quantization for Values (meaning that all configurations are the same as in KVQuant, except for the datatype). We compare against both uniform (int3) and non-uniform (nf3) [8] approaches, as well as with using “unweighted” k-means (i.e., not sensitivity-weighted) and “Fisher-weighted k-means” (without accounting for per-channel scaling factors) to compute the non-uniform quantization signposts. Note that there is slight variation in average bitwidth across models due to the differing hidden dimensions. Results report perplexity with 2K/4K/8K sequence length for LLaMA, Llama-2, and Mistral, respectively.
Method LLaMA Llama-2 Mistral-7B Avg. Num. Bits
7B 13B 30B 65B 7B 13B 70B
@@ Appendix G Sensitivity-Weighted Non-Uniform Quantization Ablations
nuq3 (Fisher-weighted) 6.01 5.34 4.41 3.74 5.49 4.83 3.26 5.03 3.00-3.02
nuq3 (KVQuant) 5.94 5.32 4.34 3.68 5.39 4.82 3.23 4.98 3.00-3.02
Appendix H Per-Vector Dense-and-Sparse Quantization Ablations
Table 10 shows the performance improvements we observe when isolating a small portion of outliers and storing them in a sparse format. We provide results both with using a single per-matrix outlier threshold, as well as with applying separate outlier thresholds per-vector. In particular, we see greater improvements by employing outlier detection with a different threshold per-channel for Keys and per-token for Values. This provides additional benefits since some values which would be considered outliers for the entire matrix are not actually outliers within a particular channel (so they are not hard to quantize). It is therefore better to directly target the outliers that will skew the quantization range for a particular channel. By removing 1% of outliers using per-vector thresholds, we can achieve an additional 0.19 reduction in perplexity for the LLaMA-7B model at 3 bits, thereby enabling 3-bit quantization with under 0.1 degradation in perplexity.
Table 10. Ablation Study: Perplexity comparison of different outlier isolation methods for LLaMA-7B on Wikitext-2. Per-vector outlier detection allows for significant accuracy improvements relative to per-tensor outlier detection. All experiments use per-token quantization for Values and per-channel quantization for Keys (pre-RoPE). “PV” refers to using per-vector outlier thresholds, and “PM” refers to using a single per-matrix outlier threshold.
@@ Appendix H Per-Vector Dense-and-Sparse Quantization Ablations
nuq3 - - 5.94 12.0
nuq3 0.1% PM 5.89 12.2
nuq3 0.1% PV 5.82 12.2
nuq3 1% PM 5.85 13.3
nuq3 1% PV 5.75 13.3
@@ Appendix I Attention Sink-Aware Quantization Ablations
Table 11 provides perplexity with and without Attention Sink-Aware quantization across different LLaMA, Llama-2, and Llama-3 models. For all bit widths (4, 3, and 2-bit) and in both dense-only and dense-and-sparse quantization settings, Attention Sink-Aware quantization consistently yields perplexity improvement. Notably, the perplexity gain is more pronounced at lower bit widths and without sparsity. We observe particularly significant improvements for Llama-3 models, with perplexity improvements of 0.25 and 0.13 PPL for nuq2-1% on Llama-3-8B and Llama-3-70B, respectively.
Table 11. Ablation Study: Perplexity with and without Attention Sink-Aware quantization on Wikitext-2 with various models. Attention Sink-Aware quantization consistently improves perplexity across different models and bit widths, particularly with lower bit widths and without sparsity.
Datatype Attention Sink-Aware LLaMA-7B LLaMA-13B Llama-2-7B Llama-2-13B Llama-3-8B Llama-3-70B
fp16 - 5.68 5.09 5.12 4.57 5.54 2.59
nuq4 X 5.72 5.14 5.17 4.62 5.64 2.65
nuq4 O 5.72 5.13 5.16 4.60 5.60 2.62
nuq4-1% X 5.69 5.10 5.13 4.59 5.57 2.60
nuq4-1% O 5.69 5.10 5.13 4.58 5.56 2.60
nuq3 X 5.94 5.32 5.39 4.82 6.10 2.99
nuq3 O 5.87 5.25 5.34 4.71 5.84 2.72
nuq3-1% X 5.75 5.15 5.18 4.62 5.67 2.66
nuq3-1% O 5.75 5.14 5.17 4.61 5.64 2.63
nuq2 X 8.47 7.29 11.20 23.34 16.63 6.79
nuq2 O 7.23 5.82 7.03 9.59 7.04 3.49
nuq2-1% X 6.05 5.39 5.47 4.85 6.29 2.95
nuq2-1% O 6.01 5.36 5.41 4.78 6.04 2.82
@@
Table 12 shows accuracy results when using offline calibration for computing the scaling factors for the Keys. For 3-bit quantization, we observe minor accuracy degradation when not employing outlier extraction methods. However, if we remove a small percentage of outliers, then the accuracy with offline calibration is the same as computing the scaling factors online per-channel during evaluation. This demonstrates that when incorporating outlier extraction methods, we are better able to perform offline calibration due to reduced sensitivity to outliers (either to outliers during calibration that exaggerate the quantization range, or to outliers during evaluation that cannot be represented accurately if there weren’t large outliers observed during calibration).
Table 13 shows the runtime for the $topk$ operation for the LLaMA-7B model (which is required for computing outlier thresholds online). It compares the runtime of the $topk$ operation with the runtime for the QKV projections, finding that the $topk$ runtime is 45% of the matrix-vector operation runtime for a single projection layer. The $topk$ operation can also be performed efficiently on the CPU, so we can actually run this operation in parallel with the subsequent linear layer matrix-vector operations on the GPU (which is possible by computing the Value projection before the Key and Query projections). This allows us to compress the activations dynamically without added runtime overhead, thereby enabling online scaling factor computation for the Value tensors.
Table 14 shows how both Fisher information computation and calibration (including k-means) per-layer take only a few minutes for the LLaMA-65B model on a typical server machine. Even if we perform calibration sequentially for each layer, the entire calibration process would take a maximum of 6 hours for the LLaMA-65B model at 4-bit precision.
Table 12. Ablation Study: Model accuracy when using offline calibration for Keys with LLaMA-7B. When incorporating outlier detection, offline calibration for Keys is able to perform comparably with online calibration. All nuq3 experiments use per-token quantization for Values and per-channel quantization for Keys (pre-RoPE), and experiments with outliers use per-vector outlier detection.
@@
nuq3 - 5.91 5.94
nuq3 1% 5.75 5.75
Table 13. $topk$ runtime on a vector of length 4096 for computing outlier thresholds when using 1% sparsity (compared with the runtime for the QKV matrix multiplications, which are 4096$\times$4096 by 4096 matrix-vector multiplications for the LLaMA-7B model). The runtime is reported on a system with an A6000 GPU and an Intel Xeon Gold 6126 CPU. We find that the runtime for the $topk$ operation is only 45% of the runtime of each matvec operation. Additionally, the $topk$ operation can be performed efficiently on the CPU; we can therefore run this operation in parallel with subsequent linear layer operations on the GPU to compress the activations dynamically without added overhead.
@@
QKV Projection GPU - 0.172
$topk$ CPU 1% 0.026
$topk$ GPU 1% 0.088
QKV Projection / $topk$ (Fused) GPU / CPU 1% 0.173
@@
For our empirical evaluation, we use 16 calibration samples of sequence length 2K from the Wikitext-2 training set (as well as the corresponding gradients) to derive the per-channel scaling factors and zero-points, and to derive the non-uniform datatypes for both Keys and Values. We measured perplexity on both Wikitext-2 and on C4 using a sequence length equal to the maximum context length of the model (2K for LLaMA, 4K for Llama-2, and 8K for Llama-3 and Mistral-7B). For baseline experiments, we use post-RoPE quantization, both since this is required from an efficiency perspective without a dedicated kernel implementation, and because it provides better accuracy when quantizing Keys per-token as shown in Appendix N.
We make several assumptions in order to estimate average bit widths and KV cache sizes for different approaches. We compute these estimates assuming a sequence length of 128K (unless otherwise specified). For integer quantization, we assume a low-precision integer offset and a 16-bit scaling factor, whereas for NormalFloat and NUQ we assume that the zero-point and offset are each 16-bit. For the sparse matrices, 32-bit integers are assumed for the per-token indices (since we need to support long sequence lengths), and the elements and per-element indices are assumed to be 16-bit. This means that for CSR, the rows are assumed to be 32-bit and the columns and values are assumed to be 16-bit, whereas for CSC, the columns are assumed to be 32-bit and the rows and values are assumed to be 16-bit.
@@
Table 15 shows perplexity across LLaMA, Llama-2, Llama-3, and Mistral models on Wikitext-2 using different datatypes. The results highlight that the NUQ datatype generally outperforms both uniformly quantized INT datatype and non-uniformly quantized NF datatype even after they are incorporated with grouping at the expense of increased bit-widths. This performance can be further improved by extracting 0.1% to 1.0% of outliers. Note that NUQ with a sparsity level 1.0% has a similar memory requirement as INT with a group size of 64, but achieves significant perplexity improvement across all models and bit widths.